# A new indeterminacy-based quantum theory

### v4

@@ 2 Main features: four conditions and two hypotheses
To simplify, throughout this paper, I will employ the familiar view that what exists are systems; a system is characterized by a collection of observables, and an observable of a system sometimes has a determinate value, where its eigenvalues represent the latter. This leads systems to "have a determinate value." Or its observables sometimes have indeterminate values, leading systems to "have an indeterminate value." Interactions are represented via QT, and some of them (which are represented via decoherence) lead systems to have a determinate value of an observable.
@@
Different ontologies can make the above view more precise and allow EnDQT to adopt a more robust realism. One may understand determinate values of systems as referring to flashes that arise or are produced under interactions, i.e., an ontology of local events in spacetime (but differently from spontaneous collapse theories and with a different interpretation of the quantum state), but there are other ways. [^14] The flash ontology was first proposed by Bell (2004) and named by Tumulka (2006). We could also view observables as representing determinables (e.g., position, energy, etc.) and determinate values as representing determinates of those determinables. Interactions give rise to a determinable with a determinate. Systems are collections of determinables, which at different moments of time, have determinates or not (e.g., having a spin-z with or without a determinate of spin-z) depending on their interactions like in the gappy version of quantum indeterminacy presented in Calosi & Wilson (2018). Quantum indeterminacy arises when we have a state of affairs constituted by a system lacking a determinate of a determinable. [^15] Alternatively, we could have an ontology of quantum properties, and this is the one I prefer (see Appendix B).
@@
Any observable $O$ of $S$, including the non-dynamical ones, outside of specific interactions of $S$ involving $O$, cannot have determinate values but rather have indeterminate values. [^18] The eigenstates of the non-dynamical observables, which are never observed in a superposition, are typically considered to be subject to superselection rules (see, e.g., Bartlett et al., 2007). These rules can be regarded as prohibiting the preparation of quantum states in a superposition, which are eigenstates of some observable and assume a coherent behavior. Rather than postulating these rules, decoherence in a widespread environment in spacetime might be used to explain this superselection (see, e.g., Earman, 2008; Giulini et al., 1995). This is the perspective taken here. However, one may object to this perspective, and EnDQT can be adapted to allow nondynamical observables of systems to always have determinate values, even when they aren’t interacting.
A new feature that EnDQT introduces is the determination capacity (DC), which will have the great benefit of allowing for the formulation of some conditions that establish when systems have determinate values, but without modifying fundamental equations of QT and allowing for the fulfillment of the above desiderata. Below, I will explain these conditions, which will establish what it takes for a system to have a determinate value and transmit the DC to others and how the former and the latter are related. I will call them Conservative Determination Conditions (CDCs). There are other possible conditions, but I found these the simplest ones, and they will suffice for the current purposes.
@@
CDC1) The determination capacity (DC) of system $X$ concerning system $Y$ (DC-Y) is the capacity that $X$ has while interacting with $Y$,
@@
i) to allow $Y$ to have a determinate value under this interaction with $X$ at $t$ that also leads $X$ to have a determinate value at $t$ and
@@
ii) to provide the DC to $Y$ concerning another system $Z$ (DC-Z) if and only if a) $Z$ starts interacting with $X$ while $Y$ is interacting with $X$, and b) $Y$ has a determinate value due to $X$.
@@
So, the DC propagates between systems via interactions because $Z$ can then have the DC concerning a system $K$ (DC-K), iff a) $K$ starts interacting with $Z$ while $Z$ is interacting with $Y$, and b) $Z$ has a determinate value due to $Y$, and so on for a system $L$ that interacts with $K$ while $K$ interacts with $Z$, etc.
How does the DC propagate more concretely? The DC propagates between systems via local interactions over spacetime, so interactions only involve systems that aren’t spacelike separated, where following the standard way,
@@
For a system $X$ to interact with system $Y$ from time $t$ to $t^{\prime}$, the quantum states of $X$ and $Y$ must at least be acted non-trivially by the Hamiltonian of interaction representing the local interaction between $X$ and $Y$ from $t$ to $t^{\prime}$.
@@
So, the reason I want $X$ to have DC-Y while interacting with $Y$, and to provide the DC-Z to $Y$ iff a) $Z$ starts interacting with $X$ while $Y$ is interacting with $X$ and b) $Z$ has a determinate value due to $Y$ is that, first, I want to provide clear criteria for systems to give rise to other systems having determinate values and to propagate that capacity to these systems. Without a criterion like CDC1), tracking when systems have the DC would be hard. Second, I also want clear criteria that establish that they can lose that capacity since, given the above desiderata, I want to allow for the possibility of arbitrary systems to be in a superposition for an arbitrary amount of time, even if they are interacting with other systems. So, we want these latter systems to lose that capacity. Third, I want to appeal $only$ to (local) interactions represented by QT for systems to have the DC and not some other criteria (except in the case of some special and plausible systems, more on this below), hence a). Fourth, I want such criteria to be plausible in the sense that it is in agreement with what we have in decoherence models and measurement-like situations where system $E$ (such as a measurement device) that gives rise to a system $S$ having a determinate value, before interacting with $S$, has a determinate value of some observable like we seem to have in the ready state of a measurement device, hence b). Note that system $Y$ could give rise to a measurement-like interaction when interacting with system $Z$. This condition will become clearer when we have all the conditions spelled out.
@@
Now, which interactions give rise to determinate values? Since my aim here is to be conservative, I will use decoherence to represent those interactions because physicists standardly use it to represent measurement-like interactions. I will now briefly explain decoherence and some of the assumptions I will make. I will also highlight with numbers some of the assumptions made in decoherence models, which EnDQT will later justify. Let’s consider a system $S$ in the following state,
@@
and an environmental system $E$ of $S$, constituted by many subsystems, interacting strongly with system S. For instance, $|\psi\rangle_{S}$ could be a superposition of spin-z eigenstates. Furthermore, $S$ will be interacting strongly with the many subsystems with a spin in a specific direction that constitutes $E$, i.e., the Hamiltonian of interaction dominates the systems’ evolution. [^19] See, e.g., Cucchietti et al. (2005) So, the dynamics will be driven by an interaction Hamiltonian, which governs or describes the interactions between systems that can affect specific observables ($i^{*})$). For simplicity, throughout this article, I will assume this kind of evolution of systems under interactions with their environment. [^20] This is the so-called quantum-measurement limit and is typically successful in describing many measurement-like interactions (Schlosshauer, 2007) In this case, the energy scales of the system-environment interaction are much larger than the energy scales associated with the self-Hamiltonians of the system and environment. More complex models of decoherence (see, e.g., Zurek, 2003, Zurek et al., 1993) where the system doesn’t interact strongly with the environment, and self-Hamiltonian also has some weight in the evolution of the system, may give rise to different observables with determinate values depending on the initial quantum states. For simplicity, I will not talk about these more complex cases here or analyze how, in these cases, SDCs could be formed. Now, let’s assume that $S$ locally interacts with E, where their interaction is represented via the standard von Neumann interaction at least approximately by $\left|s_{i}\right\rangle_{S}\left|E_{0}\right\rangle_{E}\rightarrow_{\widehat{U}}\left|s_{i}\right\rangle_{S}\left|E_{i}(t)\right\rangle_{E}$ for all $i$ and
@@
Under a Hamiltonian of interactions describing the interactions between the target system and many systems), and systems having randomly distributed initial states and coupling constants ($ii^{*})$), in decoherence models we obtain that $\left\langle E_{i}(t)\mid E_{l}(t)\right\rangle_{E}$ quickly decreases over time until $\left\langle E_{i}(t)\mid E_{l}(t)\right\rangle_{E}\approx 0$ when has $E$ is constituted by many systems. The recurrence time of this term (back to not being significantly small in comparison with the other terms) in this case tends to be so large that it can exceed the universe’s age, giving rise to a quasi-irreversible process mathematically speaking. Note that I will provide a distinct and "more pragmatic" sense of irreversibility below concerning other features of the environment that should be distinguished from this sense of irreversibility.
So, when this particular quasi-irreversibility occurs for a system $S$ interacting with $E$ being modeled in the above way, I will consider that $S$ was decohered by system $E$ or the above states of $S$ (also often called pointer states) were decohered by the states $\left|E_{i}(t)\right\rangle_{E}$ of $E$ or by $E$.
Importantly, note that decoherence here does not necessarily mean the process of destruction of interference but whatever is represented via these models. As it will be clearer, it will only refer to that process when $E$ has the DC. I will call the process that is modeled via decoherence in the sense that the environmental system has the DC, fundamental decoherence. More concretely, I will assume that when $E$ has the DC, the reduced density operator $\hat{\rho}_{S}$ can be used to predict the determinate values and resultant statistics of the consequences of this interaction (i.e., the determinate values of $S$ and $E$), as well as the timescale in which we can update the state of $S$ to one of the $\left|s_{i}\right\rangle_{S}$ under decoherence iii*). Relatedly, as it will become clearer below when $E$ has the DC, this model can directly account for the disappearance of interference effects due to $S$ in situations where it interacts with $E$. So, in this case, when states of the environment become extremely distinguishable under interactions between $S$ and $E$ over time, we have,
@@
So, given the above, I will assume that
CSC2) Interactions between system $X$ and a set of systems that form a larger system $Y$, which have the DC, lead system $X$ to have a certain determinate value, where the distinguishability of the physical state of $Y$ at $t$ concerning the possible determinate values of $X$ allows us to infer if $X$ will have a determinate value among the possible ones and when that happens. Such distinguishability is inferred via the (fundamental) decoherence of $X$ by $Y$ at $t$, and where it’s indeterministic which values will arise.
@@
In the simple situations that we will be concerned with here where the Hamiltonian of interaction dominates the evolution of the system, in order for system $Y$ to have a determinate value $v$ of $O$, i) the observable $O$ of $Y$ that is monitored by system $X$ that has the DC, and whose eigenstates are decohered by $X$ in the sense above, has to at least approximately commute with the Hamiltonian of interaction representing the interaction between $X$ and $Y$ (commutativity criterion), [^21] See Schlosshauer (2007). In more complex models of decoherence (see previous footnote), note that this monitoring may be indirect, such as the decoherence of momentum in more complex models of decoherence than the ones mentioned here (Zurek et al., 1993), where there is direct monitoring of the position. The latter is contained in the Hamiltonian of interaction of the system (but not the former), and that’s why it is considered that the decoherence of the momentum is indirect. and ii) where the eigenvalues of $O$ include $v$.
@@
Thus, the determinate values that arise are the ones that are dynamically robust under interactions with certain systems that have the DC. Note that time instants such as $t^{\prime}$ above or time intervals around $t^{\prime}$, where a system has a determinate value due to interactions with environmental systems, from now on will be represented and inferred via the time that the overlap terms above go quasi-irreversibly to zero under decoherence due to a system having the DC (i.e., the decoherence timescale). So, the above overlap terms going quasi-irreversibly to zero will allow us to infer if the local interactions between $S$ and the environmental systems that have the DC, succeeded in giving give to $S$ having a determinate value of $\mathrm{O}$. This success is inferred to occur at $t^{\prime}$ or around $t^{\prime}$. Furthermore, it’s indeterministic which value $S$ will have among the possible ones, where the latter is given by the eigenvalues of $O$ [^22] For simplicity, I will not address here the case where we don’t have maximum distinguishability of the states of the environment $E$, concerning the states of the target system. Roughly, this case can be inferred by the overlap terms of the environment not being zero or one stably over time. In these cases, the target system won’t have a determinate value due to $E$. See Pipa (2024) for an ontology of quantum properties that addresses this case.
@@
The reader familiar with decoherence might have found the above description of fundamental decoherence as missing a sometimes cited ingredient of decoherence, which is that decoherence is associated with the openness of the environment or is associated with "the entanglement of the degrees of freedom of the system" with inaccessible or uncontrollable environments that make the state of the whole larger system hard/impossible to reverse. I will now relate this feature to fundamental decoherence models. First, let’s call the models of decoherence where we don’t know if the environmental systems have the DC or that is not specified, pragmatic decoherence. In contrast with fundamental decoherence models, pragmatic decoherence models involve other considerations that go beyond the model itself such as if the environment is open or not. Pragmatic irreversible decoherence models are models that involve situations where it’s considered that is very difficult or impossible to reverse the process represented by them because they concern open environments (which we lack access to) and/or the interaction with many systems that are difficult to control. The processes represented by the pragmatic irreversible decoherence models are the processes that we normally call decoherence. However, along with my explanation of the other CDCs, I will explain below that the processes involving fundamental decoherence resemble, in important ways, the processes represented by the pragmatic irreversible decoherence models. Afterward, I will put the relation between these models in a clearer foundation.
Returning to the CSCs, we can now use CSC2) to spell out CSC1) in terms of fundamental decoherence.
@@
i) to decohere $Y$ at $t$, which leads both systems to have a determinate value at $t$. Let’s suppose that system $S$ in eq.(2) is an instance of $X$, and system $E$ is an instance $Y$. The possible values of $X$ would be represented by the eigenvalues of the observable that the quantum states $\left|s_{i}\right\rangle_{S}$ of $S$ in eq.(2) are eigenstates of. The possible values of $Y$ would be represented by the eigenvalues of the observable that the quantum states $\left|E_{i}(t)\right\rangle_{E}$ in eq.(2) are eigenstates of and
ii) to provide the DC-Z to $Y$ if and only if ii-a) $Z$ starts interacting with $Y$ while $Y$ is interacting with $X$ and ii-b) $Y$ is decohered by $X$.

### v5

@@ 2.1 The four conditions
Before presenting the four conditions concerning how the capacity to have determinate values, and to provide other systems that capacity, spreads through interactions, I will explain some background key assumptions that I will make. To simplify, throughout this article, I will employ the familiar view that what exists are systems; a system is characterized by a collection of observables, and an observable of a system sometimes has a determinate value, where its eigenvalues represent the latter. This leads systems to “have a determinate value.” Or its observables sometimes have indeterminate values, leading systems to “have an indeterminate value.” Interactions are represented via QT, and some of them (which are represented via decoherence) lead systems to have a determinate value of an observable.
@@
Different ontologies can make the above view more precise and allow EnDQT to adopt a more robust realism. One may understand determinate values of systems as referring to flashes that arise or are produced under interactions, i.e., an ontology of local events in spacetime (but differently from spontaneous collapse theories and with a different interpretation of the quantum state), but there are other ways. [^18] The flash ontology was first proposed by [12] and named by [91]. We could also view observables as representing determinables (e.g., position, energy, etc.) and determinate values as representing determinates of those determinables. Interactions give rise to a determinable with a determinate. Systems are collections of determinables, which at different moments of time, have determinates or not (e.g., having a spin-z with or without a determinate of spin-z) depending on their interactions like in the gappy version of quantum indeterminacy presented in [23]. Quantum indeterminacy arises when we have a state of affairs constituted by a system lacking a determinate of a determinable. [^19] Alternatively, we could have an ontology of quantum properties, and this is the one I prefer. See Appendix C and [81]. To be clear, according to EnDQT, determinacy and indeterminacy are objective features of the world. EnDQT postulates that (like other theories postulate other things). Determinacy is also associated with classicality (i.e., phenomena associated with classical physics) and measurement records, and our evidence for determinacy will involve these phenomena, as we will see.
@@
Any observable $O$ of $S$, including the non-dynamical ones, outside of specific interactions of $S$ involving $O$, cannot have determinate values but rather have indeterminate values.
The assumption that systems have indeterminate values of any observable by default except during certain interactions comes from the perspective that the so-called links that connect state assignments to assignments of determinate values are idealizations that never occur in practice. For instance, [96] provides good arguments for why the so-called Eigenstate-Eigenvalue Link is not a very useful assumption. As we have seen in the introduction, this link says that a system has a determinate value of an observable $O$ if and only if the system is in a state that is an eigenstate of $O$. A system rarely is in an eigenstate of such observable. When it is and if it is (since this occurs upon typically idealized projective measurements), it is, for a very brief moment of time, quickly evolving out of those states. Decoherence and decohering interactions provide a much less problematic way of assigning determinate values to the observables of systems because it is less grounded in idealized assumptions. [^22] Furthermore, as argued by [43], the so-called environmentally-induced superselection can be used to justify why systems are never found in a superposition of the eigenstates of certain observables, always having determinate values of that observable (e.g., mass or electric charge) upon interactions. More concretely, the eigenstates of the so-called non-dynamical observables, which are observables that are never observed in a superposition, are typically considered to be subject to superselection rules (see, e.g., [8]). These rules can be regarded as prohibiting the preparation of quantum states in a superposition, which are eigenstates of some observable (e.g., electric charge) and assume a coherent behavior. Rather than postulating these rules, decoherence in a widespread environment in spacetime might be used to explain this superselection (see, e.g., [36, 43]). EnDQT will assume this perspective here (not that if systems evolve back into a superposition of such initially superselected states, if they are always subject to those interactions, they will very likely maintain the original eigenstate). However, the validity of EnDQT is not tied to this assumption. Instead, for example, we can assume that there are certain non-dynamical observables that always have determinate values. Or, we can rather assume that these non-dynamical observables should be regarded more like constants and considered to be less fundamental than the dynamical ones. It is rather the behavior of systems due to dynamical observables (which are subject to decoherence) that is more fundamental and explains the features of non-dynamical observables, where this behavior is described/governed by laws. An important feature that EnDQT introduces is, as I have mentioned, the Determination Capacity (DC). The DC will have the great benefit of allowing for the formulation of a set of conditions that establish when systems have determinate values, but without modifying fundamental equations of QT and allowing for the fulfillment of the desiderata explained in Section 1.
Below, I will explain the conditions that will establish how the DC is transmitted between systems, where the DC is the capacity to give rise to other systems having a determinate value via interactions and to transmit the DC to other systems in the non-relativistic regime. Also, these conditions establish what it takes for a system to have a determinate value in this regime. [^23] In a relativistic quantum field theoretic of these conditions, some further precisifications will be needed, but they won’t affect their core features [82]. Due to their conservativeness, I will call them Conservative Determination Conditions (CDCs). CDCs are laws or law-like features that describe/govern how determinate values arise and the capacity to have determinate values is transmitted via interactions. There are alternative determination conditions, but I found these to be the simplest and most conservative ones, as well as the ones that more appropriately describe a chain of decohering local interactions in the non-relativistic regime.
@@
CDC1) For EnDQT, the world is fundamentally constituted by systems with indeterminate values of any observable. The determination capacity (DC) of system $X$ concerning system $Y$ (DC-Y) is the capacity that $X$ has while interacting locally with $Y$,
@@
i) to allow $Y$ to have a determinate value under the interaction with $X$ that also leads $X$ to have a determinate value, where $X$ and $Y$ have a determinate value in the same spacetime region,
@@
ii) to provide the DC to $Y$ concerning another system $Z$ (DC-Z) if and only if a) $Z$ starts interacting locally with $Y$ while $Y$ is already interacting with $X$, and b) $Y$ has a determinate value due to $X$ and $Z$ doesn’t disturb this process.
@@
So, the DC propagates between systems via interactions because $Z$ can then have the DC concerning a system $K$ (DC-K) if and only if a) $K$ starts interacting with $Z$ while $Z$ is already interacting with $Y$, and b) $Z$ has a determinate value due to $Y$, and so on for a system $L$ that interacts with $K$ while $K$ [^24] Note that more systems may be needed so that $L$ has a determinate value. interacts with $Z$, etc. Note that $X$ having a determinate value and $Y$ having a determinate value in the interaction in i) is the same event (i.e., it occurs in the same spacetime region. This is why, for EnDQT, interactions give rise to determinate values.
As we can see, CDC1) doesn’t invoke the EEL to assign determinate values to observables of quantum systems. Rather, it considers that systems only have determinate values of observables during interactions with certain other systems. Otherwise, systems have indeterminate values of any observable. As we will see, the motivations for these features of CDC1) are connected with the aim of EnDQT of fulfilling the desideratum LC) of maintaining locality. Note that the assumption that systems only have determinate values under interactions doesn’t mean that these values are less objective. The analogy here is with the binding energy of atoms. Binding energy arises under interactions, but we tend to consider binding energy as existing out there.
I consider the assumption that the target system and environment have a determinate value in the same spacetime region, i.e., while interacting, as a conservative one. As foreshadowed in the previous section, I will use decoherence to establish when systems have determinate values. Since decoherence involves entanglement between the system and its environment, this assumption agrees with how systems become entangled at a time in local interaction. If system $A$ gets entangled with system $B$, system $B$ gets entangled with $A$. The quantum framework doesn’t distinguish which one happens first. This assumption is also important in order for EnDQT to fulfill its aims. If these systems had determinate values at the same time in different spatial regions, the related events would be space-like separated, and according to special relativity, we would have relativistic reference frames where one event happens first, then the other, and thus we would fall into relationalism since which system has a determinate value would vary with the reference frame unless we favored a preferred reference frame. Furthermore, if this were the case, it would be more problematic to establish that the environment gave rise to the target systems having a determinate value.
Now, why not consider instead that the events involving elementary systems of the environment having determinate values are time-like separated (i.e., occurring at different times in the same spatial region) rather than occurring in the same spacetime region? The reason why is more obvious but worth mentioning. This alternative assumption would be unsatisfactory since it could give rise to a certain future dependency in terms of when determinate values arise. To see this, let’s assume that each system of the environment has the DC and interacts with the target system, having a determinate value in this interaction before the target system, and thus the target system doesn’t yet have a determinate value (so the events where they have determinate values are time-like separated). Furthermore, let’s consider that we have a process that doesn’t give rise to decoherence because the environmental overlap terms oscillate between zero and non-zero. So, the entanglement between the system and the target system would oscillate over time. Since the environmental systems would already have determinate values, due to their entanglement with the target system (at least for some time), this would allow for an observer to gain knowledge about the determinate value of the target system by looking at the environmental systems that are highly entangled with the target system at least for a brief moment in time. Hence, we would have a contradiction with the assumption that the target system doesn’t yet have a determinate value. Although these values would oscillate quickly, this would still be a contradiction. To deal with this situation, one could argue that systems of the environment would only have determinate values in the present if there was decoherence due to these systems in the future. However, this would give rise to future dependency and likely either superdeterminism or retrocausality. Hence, I have considered that the target system and the environmental systems have a determinate value in the same spacetime region while they interact. [^25] One may object to the assumption that decohering systems of the environment and the target system have determinate values in the same spacetime region by mentioning the collisional models of decoherence (see [108, 89] and references therein). In these models, one might be tempted to consider that the collisions that give rise to decoherence happen over time, and hence, the events that give rise to decoherence should be time-like separated. However, note that if collisions involving many systems happened at the same time, there wouldn’t exist a significant change in these models. Also, the fundamental theory of the world will likely be one of quantum fields, which doesn’t use the variable position, and ultimately, these models are only effective. The assumption regarding systems having determinate values in the same spacetime region shouldn’t be strange. In physics, we are used to reify features that arise due to interactions, such as the bonding energy of atoms.
Given CDC1), the DC is transmitted between systems via local interactions over spacetime,
@@
For a system $X$ to interact with system $Y$ from time $t$ to $t^{\prime}$, the quantum states of $X$ and $Y$ must at least evolve from $t$ to $t^{\prime}$ under the Hamiltonian of interaction representing the local interaction between them. [^26] So, the Hamiltonian of interaction should represent interactions in the same spacetime region. This assumption can be made more precise via quantum field theory.
@@
As I have mentioned in the previous section, the reasons for posing CDC1-ii) are that, first, I want to provide clear criteria for systems to give rise to other systems having determinate values and to transmit that capacity to these systems. Without a criterion like CDC1), tracking when systems have the DC would be hard. Second, I also want clear criteria that establish that they can lose that capacity since, given the above desiderata, I want to allow for the possibility of arbitrary systems to be in a superposition for an arbitrary amount of time, even if they are interacting with other systems. So, we want these latter systems to lose that capacity. Third, I want to appeal $only$ to (local) interactions represented by QT for systems to have the DC and not some other criteria (except in the case of some special and plausible systems; more on this below). Hence, I have also assumed CDC-ii-a). [^27] Note that establishing that interactions are local would add some complications in the quantum field theoretic case because, in that domain, we will want the quantum field systems that transmit the DC to be localized in a single bounded spacetime region. Fourth, I want such criteria to be plausible in the sense that it is in agreement with what seems to occur in decoherence models and measurement-like situations where system $E$ (such as a measurement device) that gives rise to a system $S$ having a determinate value, before interacting with $S$, has a determinate value of some observable like we seem to have in the ready state of a measurement device (i.e. before the measurement device does its measurement). Hence, I have assumed CDC1-ii-b). Now, which interactions give rise to determinate values? As I have explained in the previous section, since my aim here is to be conservative, I will use decoherence to represent those interactions because physicists standardly use it to represent measurement-like interactions.
@@
I will now explain briefly decoherence and some of the assumptions I will make when interpreting it. I will also highlight with numbers 1)1)1 ) and 2)2)2 ) two of the assumptions made in decoherence models, which EnDQT will later justify. Let’s consider a system $S$ in the following state with complex non-zero coefficients $\alpha_{i}$,
@@
and an environmental system $E$ of $S$, constituted by many subsystems, interacting strongly with system $S$. For instance, $|\psi\rangle_{S}$ could be a superposition of spin-z eigenstates. Furthermore, $S$ will be interacting strongly with the many subsystems with a spin in a specific direction that constitutes $E$, i.e., the Hamiltonian of interaction $H_{int}$ dominates the systems’ evolution. [^28] I.e., the spectral frequencies available in $H_{int}$ are much higher than the ones of the self-Hamiltonian of the systems. See, e.g., [29]. So, 1)1)1 ) the dynamics will be driven by an interaction Hamiltonian, which governs or describes the interactions between systems that can affect specific observables. For simplicity, throughout this article, I will assume this kind of evolution of systems under interactions with their environment. [^29] This is the so-called quantum-measurement limit and is typically successful in describing many measurement-like interactions ([89]) In this case, the energy scales of the system-environment interaction are much larger than the energy scales associated with the self-Hamiltonians of the system and environment. More complex models of decoherence (see, e.g., [110]) where the system doesn’t interact strongly with the environment, and self-Hamiltonian also has some weight in the evolution of the system may give rise to different observables with determinate values depending on the initial quantum states. For simplicity, I will not talk about these more complex cases here or analyze how, in these cases, SDCs could be formed. Now, let’s assume that $S$ locally interacts with $E$, where their interaction is represented via the standard von Neumann interaction at least approximately by
$\left|s_{i}\right\rangle_{S}\left|E_{0}\right\rangle_{E}\rightarrow_{\widehat{U}}\left|s_{i}\right\rangle_{S}\left|E_{i}(t)\right\rangle_{E}$ (4)
for all $i$ or
@@
Under the Hamiltonian of interaction describing the interactions between the target system and many systems, and 2)2)2 ) systems having randomly distributed initial states and coupling constants, [^30] A nonrandom initial distribution of the quantum states of the environment can lead to not having so many phases that would cancel each other out in the off-diagonal terms of the density matrix, not, therefore, leading to decoherence. in decoherence models we obtain that $\left\langle E_{i}(t)\mid E_{l}(t)\right\rangle_{E}$ quickly decreases over time until $\left\langle E_{i}(t)\mid E_{l}(t)\right\rangle_{E}\approx 0$ when has $E$ is constituted by many systems. The recurrence time of this term (back to not being significantly small in comparison with the other terms) in this case tends to be so large that it can exceed the universe’s age, giving rise to a quasi-irreversible process, mathematically speaking. Note that I will provide a distinct and “more pragmatic” sense of irreversibility below concerning other features of the environment that should be distinguished from this sense of irreversibility.
So, when either eq. (4) or (5), together with (6) hold to describe a system $S$ interacting with $E$, I will consider that $S$ was decohered by $E$ or that the superposition of the states of $S$ (also often called pointer states) were decohered by the states $\left|E_{i}(t)\right\rangle_{E}$ of $E$. The observable of $S$ formed by these pointer states is the pointer observable. [^31] Although the processes represented by eq. (4) are not what is often called decoherence, I will call them that since (more realistically) they may hold only approximately. I consider these processes rare since, typically, systems aren’t in an eigenstate of some observable. The latter should rather be seen as an idealization. See, e.g., [96].
Importantly, note that by decoherence here, it does not necessarily mean the process of destruction of interference (like it is often assumed to mean) but rather whatever is represented via these models. As it will be clearer, it will only refer to the process of destruction of interference when $E$ has the DC. I will call the process that is modeled via decoherence that involves an environmental system having the DC, fundamental decoherence. More concretely, I will assume that when $E$ has the DC, the reduced density operator $\hat{\rho}_{S}$ that is obtained from eqs. (4) or (5) can be used to predict the determinate values and resultant statistics of the consequences of this interaction (i.e., the determinate values of $S$ and $E$), as well as the timescale in which we can update the state of $S$ to one of the $\left|s_{i}\right\rangle_{S}$ under decoherence. Relatedly, as it will become clearer below when $E$ has the DC, this model can directly account for the disappearance of interference effects due to $S$ in situations where it interacts with $E$. So, in this case, when states of the environment become extremely distinguishable under interactions between $S$ and $E$ over time, we have,
@@
So, given the reasons above, I will assume that
CDC2) Interactions between system $X$ and a set of systems that form a system $Y$ that may be larger, which has the DC, lead system $X$ to have a certain determinate value, where the distinguishability of the physical state of $Y$ concerning the possible determinate values of $X$ allows us to infer if $X$ will have a determinate value among the possible ones and when that happens. Such distinguishability is inferred via the (fundamental) decoherence of $X$ by $Y$, and where it is indeterministic which one of the values will arise among the possible ones.
Thus, decoherence for EnDQT is a tool to infer how and when determinate values arise, where such values arise via an indeterministic process. More on this below.
@@
In the simple situations that we will be concerned with here where the Hamiltonian of interaction dominates the evolution of the system, in order for system $Y$ to have a determinate value $v$ of $O_{Y}$, i) the observable $O_{Y}$ of $Y$ that is monitored by system $X$ that has the DC, and whose eigenstates are decohered by $X$ in the sense above, has to at least approximately commute with the Hamiltonian of interaction $H_{XY}$ representing the interaction between $X$ and $Y$, i.e., $[H_{XY},O_{Y}]\approx 0$ (this is the so-called commutativity criterion), [^32] See [89] and references therein. Note that $[H_{XY},O_{Y}]\approx 0\Rightarrow[H_{XY},O_{Y}\otimes\mathbb{I}_{Y}]\approx 0$. Furthermore, this criterion implies that all terms in a Hamiltonian of interaction $H_{int}$ will individually satisfy this criterion. In more complex models of decoherence (see previous footnote), note that this monitoring may be indirect, such as the decoherence of momentum in more complex models of decoherence than the ones mentioned here [110], where there is direct monitoring of the position. The latter is contained in the Hamiltonian of interaction of the system (but not the former), and that’s why it is considered that the decoherence of the momentum is indirect. and ii) where the eigenvalues of $O_{Y}$ include $v$.
@@
Thus, the determinate values that arise are the ones that are dynamically robust under interactions with certain systems that have the DC. Note that the times where a system has a determinate value due to interactions with environmental systems will be represented and inferred via the time that the overlap terms above go quasi-irreversibly to zero under decoherence due to a system having the DC (i.e., the decoherence timescale). So, the above overlap terms going quasi-irreversibly to zero will allow us to infer if the local interactions between $S$ and the environmental systems that have the DC succeeded in giving give to $S$ having a determinate value of $\mathrm{O}$. Furthermore, it is indeterministic which value $S$ will have among the possible ones, where the latter is given by the eigenvalues of $O$ [^33] For simplicity, I will not address here the case where we don’t have maximum distinguishability of the states of the environment $E$, concerning the states of the target system. Roughly, this case can be inferred by the overlap terms of the environment not being zero or one stably over time. Furthermore, in these situations, the target system won’t have a determinate value due to $E$. See Appendix C and [81] for an ontology of quantum properties that addresses this case.
@@
The reader familiar with decoherence might have found the above description of fundamental decoherence as missing an often-cited ingredient of decoherence, which is that decoherence is associated with the openness of the environment or is associated with “the entanglement of the degrees of freedom of the system” with inaccessible or uncontrollable environments that make the state of the whole larger system hard/impossible to reverse. I will now relate this feature to fundamental decoherence models. First, let’s call the models of decoherence where we don’t know if the environmental systems have the DC or that is not specified, pragmatic decoherence models. In contrast with fundamental decoherence models, pragmatic decoherence models involve other considerations that go beyond the model itself, such as whether the environment is open or not. Pragmatic irreversible decoherence models are models that involve situations where it’s assumed that it is very difficult or impossible to reverse the process represented by them because they concern the interaction with many systems that are difficult to control and which often involve open environments. Importantly, these are the assumptions that go beyond the models of decoherence themselves, being rather extra assumptions postulated by the modeler. The processes represented by the pragmatic irreversible decoherence models are the processes that we normally call decoherence. However, along with my explanation of the other CDCs, I will explain below that the processes involving fundamental decoherence resemble, in important ways, the processes represented by the pragmatic irreversible decoherence models. Afterward, I will put the relation between these models in a clearer foundation.
Returning to the CDCs, we can now use CDC2) to spell out CDC1) in terms of fundamental decoherence.
@@
i*) to decohere $Y$, which leads both systems to have a determinate value of an observable O during their interaction, where, in the absence of these kinds of interactions, systems have indeterminate values of any observable. Let’s suppose that system $S$ in eq.(5) is an instance of $X$, and system $E$ is an instance of $Y$. The possible values of $X$ are represented by the eigenvalues of the observable that the quantum states $\left|s_{i}\right\rangle_{S}$ of $S$ in eq.(5) are eigenstates of. The possible values of $Y$ are represented by the eigenvalues of the observable that the quantum states $\left|E_{i}(t)\right\rangle_{E}$ in eq.(5) are eigenstates of and
ii*) to provide the DC-Z to $Y$ if and only if ii-a) $Z$ starts interacting with $Y$ while $Y$ is interacting with $X$ and ii-b) $Y$ is decohered by $X$, and $Z$ doesn’t disturb this process, i.e., driving away to other states, the states of $Y$ that are being decohered by $X$.
@@
I want to emphasize that the CDSs are just one possible condition, and there are other possible ones. As we can see, via the CDC1*) and CDC1), I postulate that in order for systems to have the DC and transmit it to other systems, they need to form a tight chain of interaction with some rules. As we will see further below, the idea is that I want to explain how sometimes we manage to isolate systems and give rise to interference phenomena via them. Imagine that this tight chain wasn’t postulated, and I had more loose determination conditions. For instance, imagine that these conditions allowed a system to simply retain the DC forever when it interacts with another system that has the DC. We would have $stray$ systems with the DC, and it becomes hard to justify why we sometimes can isolate systems from an SDC and, thus, from systems with the DC.